OpenStudy (anonymous):

Given n points was incircle have radius 1. Two points called goog pear if the points have distance more than sqrt(2). Let the circle have m good pear, prove that 3m<=n^2

hero (hero):

You should draw a picture of the problem. That would help.

OpenStudy (anonymous):

OK. my picture was on... but i dont have idea to solve this problem

OpenStudy (anonymous):

This is just Turan's theorem: there are no four-cliques, so the best we can do is to divide the points into three as-equal-as-possible groups.

OpenStudy (anonymous):

Define a graph whose vertex set is the \(n\) points, with an edge connecting two points if and only if they have distance larger than \(\sqrt{2}\) from each other. This graph contains no clique with four vertices...because in the best condition u that for will lie on the perimeter of circle... so by Turan's theorem maximum number of edges (its equal to maximum number of good pairs) is \[\frac{r-1}{r} \frac{n^2}{2}\]because we know that Among the n-vertex simple graphs with no (r + 1)-cliques, Turan's graph has the maximum number of edges. here we have \(r=3\) so \[m \le\frac{r-1}{r} \frac{n^2}{2}\]\[m \le\frac{3-1}{3} \frac{n^2}{2}=\frac{n^2}{3}\]\[3m \le n^2\]